OpenStudy (anonymous):

solving radical expressions and rational exponents. (125^2/3)^1/2

OpenStudy (anonymous):

\[(125^{\frac{2}{3}})^{\frac{1}{2}}\] \[=125^{\frac{2}{3}\times \frac{1}{2}}\] \[=125^{\frac{1}{3}}\] raise to the power of one third means take the cube root so answer is 5

OpenStudy (anonymous):

okay. im new to this so i dont know if you will be able to see my reply. but thanks.